# Dynamics and modulation of cosmic ray modified magnetosonic waves in a galactic gaseous rotating plasma

### v2

@@ IV.1 KdVB Equation Derivation
In this section, we are interested in the non-linear evolution of progressive shocks of the one-dimensional plane in a frame that moves along the $x$-axis with the wave phase velocity. For this purpose, we derive an evolution equation for small-amplitude shocks and study their characteristics in physical parameter space. To employ the standard reductive perturbation technique, we stretch the space and time coordinates,
@@
$v_{z}=0+\epsilon v_{y1}+\epsilon^{2}v_{y2}+\dots,$
@@
. Furthermore, we assume that $\eta\sim\eta_{0}\epsilon^{1/2}$ and $\kappa\sim\kappa_{0}\epsilon^{1/2}$ to make the perturbation evolution equation consistent, where $\eta_{0}$ and $\kappa_{0}$ are of the order of unity. The parameters $\eta$ and $\kappa$ are expanded in the lowest order of $\epsilon$ to take into account their effects for the propagation of dissipative magnetosonic shock waves numerically, which are valid as resistive and diffusive terms in Eq. (10) and (12), respectively, are almost linearly proportional to the number density. Similar assumptions have been made previously in many experimental situations nakamura1999observation ; nakamura2001observation . For large values of $\eta_{0}$ and $\kappa_{0}$, one can choose a higher order of $\epsilon$ (i.e., $\eta\sim\epsilon\eta_{0}$, $\kappa\sim\epsilon\kappa_{0}$) and in those cases sharp rising shock waves are noticed compared to oscillatory and monotonic shock waves (for $\eta\sim\epsilon^{1/2}\eta_{0}$, $\kappa\sim\epsilon^{1/2}\kappa_{0}$).
@@
$v_{x1}=\lambda\rho_{1}=\lambda B_{1},P_{g1}=C_{g}^{2}B_{1},v_{z1}=\lambda\cot\theta B_{1}$
@@
Similarly, we obtain a system of equations in the next order of $\epsilon$. Eliminating those perturbed quantities from the system and using the results given in Eq. (IV.1), we derive the following KdVB equation,
$\frac{\partial B}{\partial\tau}+QB\frac{\partial B}{\partial\xi}+R\frac{\partial^{3}B}{\partial\xi^{2}}=S\frac{\partial^{2}B}{\partial\xi^{2}},$ (21)
@@ IV.2 Soliton Solutions
If the effects of cosmic ray diffusivity ($\kappa_{0}$) and magnetic resistivity ($\eta_{0}$) become negligible, one can obtain the following KdV equation, disregarding their effects from Eq. (21), as
$\frac{\partial B}{\partial\tau}+QB\frac{\partial B}{\partial\xi}+R\frac{\partial^{3}B}{\partial\xi^{2}}=0,$ (25)
@@
$\mathrm{sech}\,{x}$ Now, we describe the transmission and interactions of two-soliton solutions for the KdV equation employing Hirota’s bilinear methodhirota1971exact . Recalling the variables $\xi$ by $\tilde{\xi}B^{1/3}$, $B$ by $6\tilde{B}{Q^{*}}^{-1}R^{1/3}$ and $\tau$ by $\tilde{\tau}$, we rewrite the KdV equation as
$\frac{\partial\tilde{B}}{\partial\tilde{\tau}}-6\tilde{B}\frac{\partial\tilde{B}}{\partial\tilde{\xi}}+\frac{\partial^{3}\tilde{B}}{\partial\tilde{\xi^{2}}}=0.$ (27)
@@
where $f=1+\epsilon\left(e^{\tilde{\theta}_{1}}+e^{\tilde{\theta}_{2}}\right)+\epsilon^{2}f_{2}$, $\tilde{\theta}_{i}=k_{i}\tilde{\xi}+\tilde{\omega}_{i}\tilde{\tau}$ ($i=1,2$) and $D$ represents Hirota’s operator. Collecting various powers of $\epsilon$ from Eq. (28), we obtain $\omega_{i}=-k_{i}^{3}$ and $f_{2}=a_{12}e^{\tilde{\theta}_{1}+\tilde{\theta}_{2}}$, with phase shift $a_{12}=\frac{(k_{1}-k_{2})^{2}}{(k_{1}+k_{2})^{2}}$. Finally, we have the two-soliton solution of Eq. (25) obtained as
$B=-12\frac{R^{1/3}}{Q^{*}}\frac{k_{1}^{2}e^{\theta_{1}}+k_{2}^{2}e^{\theta_{2}}+a_{12}e^{\theta_{1}+\theta_{2}}(k_{2}^{2}e^{\theta_{1}}+k_{1}^{2}e^{\theta_{2}})+2(k_{1}-k_{2})^{2}e^{\theta_{1}+\theta_{2}}}{(1+e^{\theta_{1}}+e^{\theta_{2}}+a_{12}e^{\theta_{1}+\theta_{2}})^{2}}$ (29)
@@
Numerically, we demonstrate the temporal evolution and phase shift of the two-soliton solution of the KdV equation by the interaction of two solitons. In Fig. 3, the time evaluation of two solitons interactions $B$ vs. $\xi$ have been plotted for different values of $\tau$. One can observe that two solitons with different amplitudes collide and exchange energy and phase. The shape and size of both solitons may alter during collision. Fig. 3 also depicts that for $\tau<0$, the soliton with higher amplitude stays behind the lesser one and as $\tau$ increases, the two solitons get closer. Subsequently, two solitons collide at $\tau=0$ and become one-soliton, after that, the soliton with higher amplitude moves faster than the lesser one, keeping the latter one behind. We noticed that when we changed the system’s parametric values for the two-soliton solution, the characteristics framework stayed consistent with what we saw for the one-soliton solution.
Figure 4: The magnetosonic shock profiles $B~{}\text{vs}~{}\chi$ are plotted [Eq. (38)] for different parametric values as mentioned in the legends. Here, ($\theta$, $\kappa$, $\eta$)= ($\pi/3$, 0.3, 0.03) and $(\Omega_{0},C_{g},C_{c})=(0.9,0.2,0.4)$ for the subplots (a) and (b), respectively. The other parameters remain unchanged as in Fig. 1.
Figure 5: The magnetosonic oscillatory shock profiles $B~{}\text{vs}~{}\chi$ are plotted for different values of the parameters as mentioned in the legends. The other parameters in the subplots, except the variation parameter, are fixed as $(C_{g},C_{c},\Omega_{0},\theta,\kappa_{0},\eta_{0})=(0.7,0.6,0.8,\pi/7,0.9,0.02)$.
@@ IV.3 Shock Wave Solutions
Figure 6: The instability growth rate $\Gamma$ vs. modulated wavenumber $K$ is plotted [Eq. (52)] for different values of parameters as mentioned in the legends. Here, $B_{0}=0.04$ and other parameters remain unchanged as in Fig. 2.
Figure 7: The three-dimensional first-order rogue wave profiles $|B_{1}|$ are displayed for $\Omega_{0}=0.7$ and $\Omega_{0}=0.9$ in the subplots (a) and (b), respectively. Here, $(C_{g},C_{c},\theta)=(0.1,0.7,\pi/3)$ and all other parameters remain unchanged as in Fig. 2.
@@
and the speed of the moving frame $v=\frac{6D^{2}}{25R}$. Therefore, the localized solution of the KdVB equation (21) is written as follows,
$B=\frac{3D^{2}}{25QR}\left[\mathrm{sech}\,^{2}{(\frac{S}{10R}\chi)}+2\left\{1+\tanh{(\frac{S}{10R}\chi)}\right\}\right],$ (38)
which describes a composite form of solitary and Burger shock structure. In other words, the solution (38) represents an admixture of $\mathrm{sech}\,^{2}$ and $\tanh$ type waveform. The graphical representations of such profiles have been displayed in Fig. 4 for distinct parametric values, as stated in the caption of the corresponding figure. The figures reveal that, due to the increments in the values of the parameters $\Omega_{0}$, $\theta$, $\eta_{0}$ and $\kappa_{0}$, the amplitude is shifted to higher values, while the width becomes narrower for the monotonic shock wave pulses. Thus, cosmic ray diffusivity and magnetic resistivity play a key role in determining the nature and magnitude of the shock pulse in a dissipative magneto-rotating plasma under the effect of Coriolis force. The effect of modified pressure by combining the cosmic ray pressure (relating parameter $C_{c}$) with thermal pressure (relating parameter $C_{g}$) has been observed from the subplot (a) of Fig. 4. It has been noticed that when $C_{c}$ is fixed, the amplitude is shifted to lower values and the width becomes wider for increasing values of $C_{g}$; however, the amplitude increases and the width becomes narrower for increasing values of $C_{c}$.
@@
which admits two fixed points, $(0,0)$ and $(2v/Q,0)$. Next, we employ the fourth-order Runge-Kutta method to solve Eq. (21) to study the effects of different parameters involved in the model, and the results obtained are displayed in Fig. 5. Subplots (a) and (b) show the effects of $C_{g}$ and $C_{c}$ on the oscillatory shock structures. We observe that as $C_{g}$ increases, the strength and width of the oscillatory shock grow, whereas they become smaller and tend to damp faster as $C_{c}$ increases. Subplots (c) and (d) illustrate how the rotational frequency $\Omega_{0}$ with the angle of rotation $\theta$ affects the oscillatory shock profiles. It is found that the oscillatory shock wave strength and width become smaller and tend to damp faster with increasing values of $\Omega_{0}$ with a fixed angle of rotation. On the other hand, the amplitude and width increase as $\theta$ increases for a fixed angle of rotation. Subplots (e) and (f) indicate that due to small increments in cosmic ray diffusivity $\kappa_{0}$ and magnetic resistivity $\eta_{0}$, the strength and width of oscillatory shock structures become smaller and tend to damp faster. The physical significance behind this phenomenon is that when the dissipative effects dominate the system, the monotone shock profile is formed, otherwise, the behaviour remains oscillatory.
@@ V MODULATIONAL INSTABILITY AND ROGUE WAVE SOLUTIONS
In this section, we are interested in studying the magnetosonic wavepackets’ modulation instability (MI) and the characteristics of rogue wave solutions. In the limit of low frequency (i.e., when the carrier wave frequency is much smaller than the plasma frequency), neglecting the effect of magnetic resistivity and cosmic ray diffusivity, we derive a nonlinear Schrödinger (NLS) equation from Eq. (21) that describes the weakly nonlinear evolution of the modulated wave packet’s envelope. To this end, we consider the solution of Eq. (21) in the form of a weakly modulated sinusoidal wave,
@@
$\begin{split}\frac{\partial}{\partial\tau}&\rightarrow\frac{\partial}{\partial\tau}-\epsilon v_{g}\frac{\partial}{\partial X}+\epsilon^{2}\frac{\partial}{\partial T},\\ \frac{\partial}{\partial\xi}&\rightarrow\frac{\partial}{\partial\xi}+\epsilon\frac{\partial}{\partial X}.\end{split}$ (42)
@@
$\begin{split}-i\omega lB_{l}^{(n)}-v_{g}\frac{\partial}{\partial T}B_{l}^{(n-1)}+\frac{\partial}{\partial T}B_{l}^{(n-2)}+A\sum_{n=1}^{\infty}\epsilon^{n}\sum_{l=-\infty}^{+\infty}\left(ilkB_{l}^{(n)}{B}_{l-l^{\prime}}^{(n-n^{\prime})}+{B}_{l-l^{\prime}}^{(n-n^{\prime}-1)}\frac{\partial}{\partial X}B_{l}^{(n)}\right)\\ +B\left(-il^{3}k^{3}B_{l}^{(n)}-3l^{2}k^{2}\frac{\partial{B}_{l}^{(n-1)}}{\partial X}+3ilk\frac{\partial^{2}{B}_{l}^{(n-2)}}{\partial X^{2}}+\frac{\partial^{3}{B}_{l}^{(n-3)}}{\partial X^{3}}\right)=0.\end{split}$ (43)
@@
The maximum growth rate is $\Gamma_{max}=N|B_{0}|^{2}$ at $K=K_{c}/\sqrt{2}$. The growth rate of instability depends on the coefficients $M$ and $N$ (which are further influenced by various plasma parameters), any changes in the plasma parameters certainly change the growth rate of instability. The coefficients $M$ and $N$ are modified due to the inclusion of cosmic ray pressure, which affects the growth of the instability. The graphs of the growth rate $\Gamma$ have been plotted against the modulated wave number $K$ in Fig. 6 for different values of rotational frequency $\Omega_{0}$, angle of rotation $\theta$ (see subplot (a)), the parameter associated with thermal pressure $C_{g}$ and cosmic ray pressure $C_{c}$ (see subplot (b)), respectively. We note that $\Gamma$ increases with the increasing value of $\Omega_{0}$ and $\theta$, however, it decreases as the values of $C_{g}$ and $C_{c}$ increase. It has been noticed that including cosmic ray pressure in the present model reduces the instability growth. Furthermore, the plots in Fig. 6 show that the growth rate increases with increasing modulated wave number $K$ until it reaches a critical growth rate (say, $\Gamma_{c}$), which varies for different plasma parameters. After reaching the critical valued growth rate, it intensely decreases for further higher values of $K$. In particular, one can find $\Gamma_{c}=0.017$ at $K=0.3$ in the subplot (a) of Fig. 6 for $k=0.4$ and $\theta=\pi/3$.
Figure 8: The first-order rogue wave pulses $|B_{1}|$ vs. $\xi$ [Eq. (53)] are plotted for different values of parameters as mentioned in the legends. Here, $K=0.4$ and all other parameters remain unchanged as in Fig. 2.
Figure 9: The second-order rogue wave pulses $|B_{2}|$ vs. $\xi$ (Eq. (54)) are plotted for different values of parameters as mentioned in the legends. Here, $K=0.4$ and all other parameters remain unchanged as in Fig. 2.

### v3

@@ IV NONLINEAR ANALYSIS
The linear analysis in Sec. III reveals that the instability or damping of the magnetosonic waves is caused by the contribution of cosmic ray diffusivity ($\kappa$) and magnetic resistivity ($\eta$). The Coriolis force and the modified pressure, due to the rotational effects and cosmic ray effects, respectively, influenced the damping rate of the wave. The damping rate is estimated as $|\gamma|\sim(C_{c}^{2}\kappa+\eta)k^{4}/2\omega_{1}^{2}$ in the linear perturbation. We investigate whether the linear perturbation can form nonlinear structures as they propagate and the nonlinear effect intervenes. Further, in the linear regime, perturbations are assumed to be very small, allowing us to disregard any nonlinear effects. In this section, we lighten the condition of small perturbation limits and intervene in nonlinear effects for which the linear theory is no longer valid, leading to the perturbation excitation of shocks, solitons, and rogue waves. To this end, we will derive evolution equations (KdVB equation in Subsection IV.1 and the NLS equation in Section V) for such nonlinear waves in the preceding sections and examine their behaviour in different parametric regimes.
@@ IV.1 KdVB Equation Derivation
In this subsection, we are interested in the non-linear evolution of progressive shocks of the one-dimensional plane in a frame that moves along the $x$-axis with the wave phase velocity. For this purpose, we derive an evolution equation for small-amplitude shocks and study their characteristics in physical parameter space. To employ the standard reductive perturbation technique, we stretch the space and time coordinates,
@@
$v_{z}=0+\epsilon v_{z1}+\epsilon^{2}v_{z2}+\dots,$
@@
Linear analysis in Sec. III reveals that wave damping occurs due to the effect of $\eta$ and $\kappa$. Thus, we assume that $\eta\sim\eta_{0}\epsilon^{1/2}$ and $\kappa\sim\kappa_{0}\epsilon^{1/2}$ to make the perturbation evolution equation consistent, where $\eta_{0}$ and $\kappa_{0}$ are of the order of unity. The parameters $\eta$ and $\kappa$ are expanded in the lowest order of $\epsilon$ to take into account their effects for the propagation of dissipative magnetosonic shock waves numerically, which are valid as resistive and diffusive terms in Eq. (10) and (12), respectively, are almost linearly proportional to the number density. Similar assumptions have been made previously in many experimental situations nakamura1999observation ; nakamura2001observation . For large values of $\eta_{0}$ and $\kappa_{0}$, one can choose a higher order of $\epsilon$ (i.e., $\eta\sim\epsilon\eta_{0}$, $\kappa\sim\epsilon\kappa_{0}$) and in those cases sharp rising shock waves are noticed compared to oscillatory and monotonic shock waves (for $\eta\sim\epsilon^{1/2}\eta_{0}$, $\kappa\sim\epsilon^{1/2}\kappa_{0}$).
@@
$v_{x1}=\lambda\rho_{1}=\lambda B_{1},P_{g1}=C_{g}^{2}B_{1},v_{z1}=\lambda\cot\theta B_{1},$
@@
We obtained this cosmic ray modified phase velocity due to the interaction of ionized gas with cosmic rays. In this expression of the phase velocity (20), if one neglects the effects of the parameters related to modification of pressure due to cosmic ray interaction, the expression agrees with the phase velocity as obtained in Ref. hager2023magnetosonic in absence of Zeeman energy. Next, we obtain a set of equations in the next order of $\epsilon$. Eliminating those perturbed quantities from the equations and using the results given in Eq. (IV.1), we derive the following KdVB equation,
$\frac{\partial B}{\partial\tau}+QB\frac{\partial B}{\partial\xi}+R\frac{\partial^{3}B}{\partial\xi^{3}}=S\frac{\partial^{2}B}{\partial\xi^{2}},$ (21)
@@ IV.2 Soliton Solutions
In the turbulence zones of ISM of spiral galaxies, such as galaxy centre, star-forming regions, and supernova remnants, generally the magnetic Reynold number is high, and the magnetic fields follow the “frozen-in-field” condition on large scales. In such regions, magnetic fields can sustain over a long period of time and reduce cosmic ray diffusive ability by trapping them for longer periods. Therefore, it is conventional to consider the cases of weak magnetic resistivity ($\eta_{0}$) and cosmic ray diffusivity ($\kappa_{0}$). Neglecting the effect of $\eta_{0}$ and $\kappa_{0}$, one can obtain the following KdV equation from Eq. (21), as
$\frac{\partial B}{\partial\tau}+QB\frac{\partial B}{\partial\xi}+R\frac{\partial^{3}B}{\partial\xi^{3}}=0,$ (25)
@@
Now, we describe the transmission and interactions of two-soliton solutions for the KdV equation employing Hirota’s bilinear methodhirota1971exact . Recalling the variables $\xi$ by $\tilde{\xi}B^{1/3}$, $B$ by $6\tilde{B}{Q^{*}}^{-1}R^{1/3}$ and $\tau$ by $\tilde{\tau}$, we rewrite the KdV equation as
$\frac{\partial\tilde{B}}{\partial\tilde{\tau}}-6\tilde{B}\frac{\partial\tilde{B}}{\partial\tilde{\xi}}+\frac{\partial^{3}\tilde{B}}{\partial\tilde{\xi^{3}}}=0.$ (27)
@@
where $f=1+\epsilon\left(e^{\tilde{\theta}_{1}}+e^{\tilde{\theta}_{2}}\right)+\epsilon^{2}f_{2}$, $\tilde{\theta}_{i}=k_{i}\tilde{\xi}+\tilde{\omega}_{i}\tilde{\tau}$ ($i=1,2$) and $D$ represents Hirota’s operator. Collecting various powers of $\epsilon$ from Eq. (28), we obtain $\tilde{\omega}_{i}=-k_{i}^{3}$ and $f_{2}=a_{12}e^{\tilde{\theta}_{1}+\tilde{\theta}_{2}}$, with phase shift $a_{12}=\frac{(k_{1}-k_{2})^{2}}{(k_{1}+k_{2})^{2}}$. Finally, we have the two-soliton solution of Eq. (25) obtained as
$B=-12\frac{R^{1/3}}{Q^{*}}\frac{k_{1}^{2}e^{\theta_{1}}+k_{2}^{2}e^{\theta_{2}}+a_{12}e^{\theta_{1}+\theta_{2}}(k_{2}^{2}e^{\theta_{1}}+k_{1}^{2}e^{\theta_{2}})+2(k_{1}-k_{2})^{2}e^{\theta_{1}+\theta_{2}}}{(1+e^{\theta_{1}}+e^{\theta_{2}}+a_{12}e^{\theta_{1}+\theta_{2}})^{2}},$ (29)
@@
Numerically, we demonstrate the temporal evolution and phase shift of the two-soliton solution of the KdV equation by the interaction of two solitons. In Fig. 3, the time evaluation of two solitons interactions $B$ vs $\xi$ have been plotted for different values of $\tau$. One can observe that two solitons with different amplitudes collide and exchange energy and phase. The shape and size of both solitons may alter during collision. Fig. 3 also depicts that for $\tau<0$, the soliton with higher amplitude stays behind the lesser one and as $\tau$ increases, the two solitons get closer. Subsequently, two solitons collide at $\tau=0$ and become one-soliton, after that, the soliton with higher amplitude moves faster than the lesser one, keeping the latter one behind. The temporal interaction of two solitons represents the mechanism for transporting energy and momentum in the ISM of spiral galaxies. Interaction between two solitons can lead to a localized amplification of energy and magnetic field. We noticed that when we changed the parametric values of the plasma model for the two-soliton solution, the characteristics framework stayed consistent with what we saw for the one-soliton solution.
Figure 4: The magnetosonic shock profiles $B~{}\text{vs}~{}\chi$ are plotted [see Eq. (38)] for different parametric values as mentioned in the legends. The amplitude is shifted to higher values and the width becomes narrower due to the increasing effect of rotational frequency ($\Omega_{0}$), angle of rotation ($\theta$), cosmic ray diffusivity ($\eta_{0}$), magnetic resistivity ($\kappa_{0}$) and the parameter $C_{c}$ related to cosmic pressure. However, the amplitude is shifted to lower values and the width becomes wider for increasing values of the parameter $C_{g}$ related to cosmic ray pressure. Here, ($\theta$, $\kappa$, $\eta$)= ($\pi/3$, 0.3, 0.03) and $(\Omega_{0},C_{g},C_{c})=(0.9,0.2,0.4)$ for the subplots (a) and (b), respectively. The other parameters remain unchanged as in Fig. 1.
Figure 5: The magnetosonic oscillatory shock profiles $B~{}\text{vs}~{}\chi$ are plotted [see Eq. (39)] for different values of the parameters as mentioned in the legends. The strength and width of the oscillatory shock grow as the parameter $C_{g}$ related to thermal pressure and the angle of rotation ($\theta$) increases, whereas they become smaller and tend to damp faster as the parameter $C_{c}$ related to cosmic pressure, rotational frequency ($\Omega_{0}$), cosmic ray diffusivity ($\kappa_{0}$) and magnetic resistivity ($\eta_{0}$) increase. The parameters in the subplots, except the variation parameter, are fixed as $(C_{g},C_{c},\Omega_{0},\theta,\kappa_{0},\eta_{0})=(0.7,0.6,0.8,\pi/7,0.9,0.02)$.
@@ IV.3 Shock Wave Solutions
Figure 6: The development of the instability growth rate $\Gamma$ [see Eq. (52)] against the modulated wavenumber $K$ is shown for different values of parameters as mentioned in the legends. As the modulated wave number $K$ rises, the growth rate increases until it hits a critical growth rate, say $\Gamma_{c}$. For progressively higher values of $K$, it sharply declines after reaching the critical value $\Gamma=\Gamma_{c}$. Here, $B_{0}=0.04$ and other parameters remain unchanged as in Fig. 2.
Figure 7: The three-dimensional first-order rogue wave profiles $|B_{1}|$ are displayed [see Eq. (53)] for two different values of the rotational frequency, $\Omega_{0}=0.7$ and $\Omega_{0}=0.9$ in the subplots (a) and (b), respectively. The increase in rotational frequency results in a decrease in the rogue wave’s amplitude and width. The observation is quite natural, as the pulse cannot gain energy from the background wave. Here, $(C_{g},C_{c},\theta)=(0.1,0.7,\pi/3)$ and all other parameters remain unchanged as in Fig. 2.
@@
and the speed of the moving frame $v=\frac{6S^{2}}{25R}$. Therefore, the localized solution of the KdVB equation (21) is written as follows,
$B=\frac{3S^{2}}{25QR}\left[\mathrm{sech}\,^{2}{(\frac{S}{10R}\chi)}+2\left\{1+\tanh{(\frac{S}{10R}\chi)}\right\}\right],$ (38)
which describes a composite form of solitary and Burger shock structure. In other words, the solution (38) represents an admixture of $\mathrm{sech}\,^{2}$ and $\tanh$ type waveform. The graphical representations of such profiles have been displayed in Fig. 4 for distinct parametric values, as stated in the caption of the corresponding figure. The figures reveal that, due to the increments in the values of the parameters $\Omega_{0}$, $\theta$, $\eta_{0}$ and $\kappa_{0}$, the amplitude is shifted to higher values, while the width becomes narrower for the monotonic shock wave pulses. Thus, cosmic ray diffusivity and magnetic resistivity play a key role in determining the nature and magnitude of the shock pulse in a dissipative magneto-rotating plasma under the effect of Coriolis force. The effect of modified pressure by combining the cosmic ray pressure (relating parameter $C_{c}$) with thermal pressure (relating parameter $C_{g}$) has been observed from the subplot (a) of Fig. 4. It has been noticed that when $C_{c}$ is fixed, the amplitude is shifted to lower values and the width becomes wider for increasing values of $C_{g}$; however, the amplitude increases and the width becomes narrower for increasing values of $C_{c}$. Due to the effect of cosmic ray modified pressure, cosmic ray diffusivity, and magnetic resistivity, the dissipation term in Eq. (21) is dominant. Therefore, Fig. 4 indicates that the energy dissipation caused by the cosmic ray effects leads to the temporal evaluation of shock wave profiles in the ISM of spiral galaxies. In other words, the ISM’s shock wave structures are key mechanisms for accelerating cosmic rays.
@@
which admits two fixed points, $(0,0)$ and $(2v/Q,0)$. Next, we employ the fourth-order Runge-Kutta method to solve Eq. (21) to study the effects of different parameters involved in the model, and the results obtained are displayed in Fig. 5. Subplots (a) and (b) show the effects of $C_{g}$ and $C_{c}$ on the oscillatory shock structures. We observe that as $C_{g}$ increases, the strength and width of the oscillatory shock grow, whereas they become smaller and tend to damp faster as $C_{c}$ increases. Subplots (c) and (d) illustrate how the rotational frequency $\Omega_{0}$ with the angle of rotation $\theta$ affects the oscillatory shock profiles. It is found that the oscillatory shock wave strength and width become smaller and tend to damp faster with increasing values of $\Omega_{0}$ with a fixed angle of rotation. On the other hand, the amplitude and width increase as $\theta$ increases for a fixed angle of rotation. Subplots (e) and (f) indicate that due to small increments in cosmic ray diffusivity $\kappa_{0}$ and magnetic resistivity $\eta_{0}$, the strength and width of oscillatory shock structures become smaller and tend to damp faster. The physical significance behind this phenomenon is that when the dissipative effects dominate the plasma environment, the monotone shock profile is formed, otherwise, the behaviour remains oscillatory.
@@ V MODULATIONAL INSTABILITY AND ROGUE WAVE SOLUTIONS
In this section, we are interested in studying the magnetosonic wavepackets’ modulation instability (MI) and the characteristics of rogue wave solutions. In the limit of low frequency (i.e., when the carrier wave frequency is much smaller than the plasma frequency), neglecting the effect of magnetic resistivity and cosmic ray diffusivity due to high Reynold number and “frozen-in-field” condition in the turbulence zones of ISM of spiral galaxies as discussed in Sec. IV.2, we derive a nonlinear Schrödinger (NLS) equation from Eq. (21) that describes the weakly nonlinear evolution of the modulated wave packet’s envelope. To this end, we consider the solution of Eq. (21) in the form of a weakly modulated sinusoidal wave,
@@
$\begin{split}\frac{\partial}{\partial\tau}&\rightarrow\frac{\partial}{\partial\tau}-\epsilon v_{g}\frac{\partial}{\partial X}+\epsilon^{2}\frac{\partial}{\partial T},\\ \frac{\partial}{\partial\xi}&\rightarrow\frac{\partial}{\partial\xi}+\epsilon\frac{\partial}{\partial X}.\\ \\ \end{split}$ (42)
@@
$\begin{split}-i\omega lB_{l}^{(n)}-v_{g}\frac{\partial}{\partial T}B_{l}^{(n-1)}+\frac{\partial}{\partial T}B_{l}^{(n-2)}+A\sum_{n=1^{\prime}}^{\infty}\sum_{l^{\prime}=-\infty}^{+\infty}\left(ilk{B}_{l-l^{\prime}}^{(n-n^{\prime})}B_{l^{\prime}}^{(n^{\prime})}+{B}_{l-l^{\prime}}^{(n-n^{\prime}-1)}\frac{\partial}{\partial X}B_{l^{\prime}}^{(n^{\prime})}\right)\\ +B\left(-il^{3}k^{3}B_{l}^{(n)}-3l^{2}k^{2}\frac{\partial{B}_{l}^{(n-1)}}{\partial X}+3ilk\frac{\partial^{2}{B}_{l}^{(n-2)}}{\partial X^{2}}+\frac{\partial^{3}{B}_{l}^{(n-3)}}{\partial X^{3}}\right)=0.\end{split}$ (43)
@@
The maximum growth rate is $\Gamma_{max}=N|B_{0}|^{2}$ at $K=K_{c}/\sqrt{2}$. The growth rate of instability depends on the coefficients $M$ and $N$ (which are further influenced by various plasma parameters), any changes in the plasma parameters certainly change the growth rate of instability. The coefficients $M$ and $N$ are modified due to the inclusion of cosmic ray pressure, which affects the growth of the instability. The graphs of the growth rate $\Gamma$ have been plotted against the modulated wave number $K$ in Fig. 6 for different values of rotational frequency $\Omega_{0}$, angle of rotation $\theta$ (see subplot (a)), the parameter associated with thermal pressure $C_{g}$ and cosmic ray pressure $C_{c}$ (see subplot (b)), respectively. We observe that $\Gamma$ increases with the increasing value of $\Omega_{0}$ and $\theta$, however, it decreases as the values of $C_{g}$ and $C_{c}$ increase. It has been noticed that including cosmic ray pressure in the present model reduces the instability growth. Furthermore, the plots in Fig. 6 show that the growth rate increases with increasing modulated wave number $K$ until it reaches a critical growth rate (say, $\Gamma_{c}$), which varies for different plasma parameters. After reaching the critical valued growth rate, it intensely decreases for further higher values of $K$. In particular, one can find $\Gamma_{c}=0.017$ at $K=0.3$ in the subplot (a) of Fig. 6 for $k=0.4$ and $\theta=\pi/3$.
Figure 8: The first-order rogue wave pulses $|B_{1}|$ vs X [see Eq. (53)] are plotted for different values of parameters as mentioned in the legends. The amplitude and width are found to be increased with the increasing value of the parameters $C_{g}$ and $C_{c}$ related to the thermal pressure and cosmic ray pressure, respectively. An increase in rotational frequency ($\Omega_{0}$) and/or angle of rotation ($\theta$) leads to the contraction of the rogue wave pulse. Here, $K=0.4$ and all other parameters remain unchanged as in Fig. 2.
Figure 9: The second-order rogue wave pulses $|B_{2}|$ vs X [see Eq. (54)] are plotted for different values of parameters as mentioned in the legends. The amplitude of the second-order pulse is higher than that of first-order rogue wave pulses as observed in Fig. 8. The effect of the parameters $\Omega_{0}$, $\theta$, $C_{g}$ and $C_{c}$ are found to be similar as that of first-order pulses. Here, $K=0.4$ and all other parameters remain unchanged as in Fig. 2.